# Integrated Radiation-Magneto-Hydrodynamic Simulations of Magnetized Burning Plasmas. I. Magnetizing Ignition-Class Designs

### v1

@@ VIII Conclusion
With the exception of the EYC design, all designs showed some degree of enhancement with applied magnet fields between 30 to 70 T, with yield enhancement between 2 to 12. This is primarily due to enhanced hotspot temperatures due to restricted electron heat conduction by the compressed magnetic fields. Alpha magnetization is also an important factor and contributes to the ignition of polar caps when significant at stagnation, e.g., N210808. These ad-hoc results address the historical assumption per Ref. jones and that igniting designs can in fact benefit from magnetization. Only for T do we start to see a non-retrievable degradation of an igniting design, where shape degradations due to magnetization are too severe for proper assembly of the implosion at stagnation. A unique observation that was not expected was that a notable 2-3 yield enhancement can be achieved by applying magnetic field strengths less than 5 T, even as low as 1 T. This observation has spurred the consideration of low-strength, potentially even permanent magnetic designs for magnetized ICF that may be able to lower the costs of deploying a pulsed-power system that currently prevent magnetized ICF experiments.
@@
An ongoing investigation into the fundamentals of magnetized ICF is underway to better understand the nature of these designs. The primary benefit of magnetization comes from a reduction in electron heat conduction perpendicular to magnetic field lines and enhanced alpha energy deposition. However, extended MHD is far more complicated than just these two effects. A follow-up manuscript will be published as Part II of this study, where we perform a cursory survey of extended MHD effects in magnetized ICF implosions. To date, reduced heat conduction and magnetized alpha deposition have been demonstrated in NIF experiments. However, the long-considered concept of magnetized mix is still ongoing and is a potentially promising reason for further pursuit of magnetized ICF.
@@
Having surveyed well-vetted and established ICF designs and their response to magnetization, ongoing work is being pursued. First, is the application of magnetization to Next Generation (NextGen) ICF designs, where we consider a future ICF facility with a 10 MJ laser driver. A NextGen facility will promise high gains likely approaching 10-50. Given experiences with robustly igniting designs like N221204 or EYC, a wiser application of magnetic fields needs to be considered. One promising avenue is to tune the magnetic field profile in addition to that of a simple axial field, e.g., a magnetic mirror design that might further restrict conduction losses.walsh3 An additional consideration is to shim the ice layer inside the capsule so that more fuel mass is accrued on the poles, taking advantage of the collimated heat flow and alpha deposition that we see in our simulations. To truly take advantage of magnetized ICF however, it will be necessary to develop new ICF designs that are specifically tailored to the effects of magnetization. A recent study performed scans of 1D capsule-only simulations with reduced conductivity and alpha magnetization coupled to a smaller set of 2D simulations. It was found that a redesign of a non-igniting model of N170601 (measured yield of 42 kJ), increasing the shell thickness and thinning the ice layer, coupled to applied magnetic fields, could result in an 18.5 enhancement in yield.walsh4 Such considerations can be extended to fully integrated hohlraum simulations of magnetized ICF implosions. Given the high dimensionality and nonlinearity of the problem, brute-force scans of parameter space would be prohibitively expensive. This suggests an opportunity for the application of numerical optimization, whether using Gaussian processes, adjoint methods, or low-cost gradient-based approaches.
@@ Appendix A Detailed Tuning Process
The first step of the ANTS tuning process used in this study involves an associated keyhole experiment, for example, keyhole N171016 for DT shot N180128, where the foot of the pulse is stretched out and the peak power is truncated so as to clearly resolve the shock velocities and timings in question using VISAR.Malone This sets the laser power multipliers before peak power. The peak power multiplier can be tuned separately since it has little impact on the VISAR data. A small suite of simulations is run where time-power multipliers are varied to sample the total parameter space with a dimensionality of 6 - 8 parameters. This suite of simulations can be done in 1D or 2D and are typically relatively fast as we do not need to resolve CBET or the late-time implosion physics such as nuclear burn. A depiction of the power multipliers used from the shock tune can be found in Fig. 18.
@@
As the last step we degrade the yield by using a buoyancy-drag model that is a function of three parameters: the mix width , which defines an initial mixing amplitude of coupled spikes and bubbles in the mix region, the premix fraction , which is the atomic fraction of ablator material (carbon or beryllium here) assumed to initially be in the DT vapor region, and the dopant mix fraction multiplier , which is an assumption with respect to the amount of tungsten or molybdenum mixed into the DT vapor and is a multiplier applied to the initial dopant fraction times the premix fraction, i.e., , where is listend in Table 1. The buoyancy-drag model is governed by an equation of the form:
@@
A caveat here is that is the burn-averaged, bith (thermal) ion temperature in the hot spot, which does not take into account Doppler shifting of the measured neutron spectrum whence the temperature is inferred, which results in a higher measured temperature than the baseline thermal . In experiments the measured is higher than that in the simulation, so our predictions should be a bit lower. Lastly, the yield for N221204 was kept high, 44% more than the experiment, to reduce the time spent tuning as well as to keep pace with recent HYBRID-E shots that had very similar designs but even higher yields (the current record at the time this study was done was N240210 at 5.2 MJ). The case examples of EYC and PSS do not have analogous experimental shots to date thus far, so best practices were assumed with respect to the power multipliers and optimal shape. PSS power multipliers were based on previous studies, e.g., Ref. maclaren2, and which inform ongoing and future experimental campaigns. EYC tuning was derived from the simulations used in Ref. maclaren1, which was originally tuned using wall opacity multipliers and careful tuning of the shape to optimize shock breakout. This simulation was redone with laser power multipliers, to be consistent with the rest of the simulations in this study, where one power multiplier was used for the foot and one for the peak.

### v2

@@ VIII Conclusion
With the exception of the EYC design, all designs showed some degree of enhancement with applied magnetic fields between 30 to 70 T, with yield enhancement between 2 to 12. This is primarily due to enhanced hotspot temperatures due to restricted electron heat conduction by the compressed magnetic fields. Alpha magnetization is also an important factor and contributes to the ignition of polar caps when significant at stagnation, e.g., N210808. Our results on specific NIF designs show that igniting systems can indeed benefit from magnetization, unlike the claim from works like Ref. Jones and Mead, 1986. Only for T do we start to see a non-retrievable degradation of an igniting design, where shape degradations due to magnetization are too severe for proper assembly of the implosion at stagnation. A unique, unexpected observation was a notable 2-3 yield enhancement can be achieved by applying magnetic field strengths less than 5 T, even as low as 1 T. This observation has spurred the consideration of low-strength, potentially even permanent-magnetic methods for magnetizing ICF targets. This could be much cheaper than the current pulsed-power method, which is not yet available for cryogenic NIF experiments.
@@
An ongoing investigation into the fundamentals of magnetized ICF is underway to better understand the nature of these designs. The primary benefit of magnetization comes from a reduction in electron heat conduction perpendicular to magnetic field lines and enhanced alpha energy deposition. However, extended MHD is far more complicated than just these two effects. A follow-up manuscript will be published as Part II of this study, where we perform a cursory survey of extended MHD effects in magnetized ICF implosions. To date, reduced heat conduction and magnetic confinement of energetic (MeV) charged fusion products have been demonstrated in NIF experiments. However, the long-considered concept of magnetized mix is still under study and is a potentially promising reason for further pursuit of magnetized ICF.
@@
Having surveyed well-vetted and established ICF designs and their response to magnetization, ongoing work is being pursued. Given experiences with robustly igniting designs like N221204 or EYC, a wiser application of magnetic fields needs to be considered. One promising avenue is to tune the magnetic field profile in addition to that of a simple axial field, e.g., a magnetic mirror design that might further restrict conduction losses.Walsh et al. (2025b) An additional consideration is to shim the ice layer inside the capsule so that more fuel mass is accrued on the poles, taking advantage of the collimated heat flow and alpha deposition that we see in our simulations. To truly take advantage of magnetized ICF however, it will be necessary to develop new ICF designs that are specifically tailored to the effects of magnetization. A recent study performed scans of 1D capsule-only simulations with reduced conductivity and alpha magnetization coupled to a smaller set of 2D simulations.Walsh et al. (2025a) It was found that a redesign of a non-igniting model of N170601 (measured yield of 42 kJ), increasing the shell thickness and thinning the ice layer, coupled to applied magnetic fields, could result in an 18.5 enhancement in yield. Such considerations can be extended to fully integrated hohlraum simulations of magnetized ICF implosions. Given the high dimensionality and nonlinearity of the problem, brute-force scans of parameter space would be prohibitively expensive. This suggests an opportunity for the application of numerical optimization, whether using Gaussian processes, adjoint methods, or low-cost gradient-based approaches.
@@ Appendix A Detailed Tuning Process
The first step of the ANTS tuning process used in this study involves an associated keyhole experiment, for example, keyhole N171016 for DT shot N180128, where the foot of the pulse is stretched out and the peak power is truncated so as to clearly resolve the shock velocities and timings in question using VISAR.Malone et al. (2004) This sets the laser power multipliers before peak power. The peak power multiplier can be tuned separately since it has little impact on the VISAR data. A small suite of simulations is run where time-power multipliers are varied to sample the total parameter space with a dimensionality of 6 - 8 parameters. A depiction of the power multipliers used from the shock tune can be found in Fig. 18.
@@
As the last step we degrade the yield by using a buoyancy-drag mix model that is a function of three parameters: the mix width , which defines an initial mixing amplitude of coupled spikes and bubbles in the mix region, the premix fraction , which is the atomic fraction of ablator material (carbon or beryllium here) assumed to initially be in the DT vapor region, and the dopant mix fraction multiplier , where the initial dopant (tungsten or molybdenum) fraction in the DT vapor is , being listed in Table 1. The buoyancy-drag model is governed by an equation of the form:
@@
A caveat here is that is the burn-averaged, birth (thermal) ion temperature in the hot spot, which does not take into account Doppler shifting of the measured neutron spectrum whence the temperature is inferred, which results in a higher measured temperature than the baseline thermal . In experiments the measured is higher than that in the simulation, so our predictions should be a bit lower. Lastly, the yield for N221204 was kept high, 44% more than the experiment, to reduce the time spent tuning as well as to keep pace with recent HYBRID-E shots that had very similar designs but even higher yields (the current record at the time this study was done was N240210 at 5.2 MJ). The case examples of EYC and PSS do not have analogous experimental shots to date thus far, so best practices were assumed with respect to the power multipliers and optimal shape. PSS power multipliers were based on previous studies, e.g., Ref. MacLaren et al., 2021, and which inform ongoing and future experimental campaigns. EYC tuning was derived from the simulations used in Ref. MacLaren et al., 2024, which was originally tuned using wall opacity multipliers and careful tuning of the shape to optimize shock breakout. This simulation was redone with laser power multipliers, to be consistent with the rest of the simulations in this study, where one power multiplier was used for the foot and one for the peak.